# New Foundations is consistent

### v22

@@ A $\beta$-support $S$ is called strong if it satisfies the additional properties that
The elements of the range of a $\beta$-specification are taken from a set of size less than $\mu$, since they are built from ingredients in $\kappa$ (support domain elements), $\beta+1$ (type indices currently in use), and $\gamma$-coding functions for $\gamma<\beta$; each of the collections from which the ingredients are taken are of size $<\mu$ and $\mu$ has cofinality at least $\kappa$. Therefore, the cardinality of the set of specifications is bounded by $\nu^{\xi}$ where $\nu<\mu$ and $\xi<\kappa$. Now, $\nu^{\xi}$ is the cardinality of the set of functions from $\xi$ to $\nu$, which is less than or equal to the cardinality of the power set of $\nu\times\xi$, which in turn is less than $\mu$ because $\mu$ is strong limit. ∎
A detail: We need to count all coding functions, allowing $\gamma<\beta$ to vary. We have $\beta<\lambda\leq{\tt cf}(\mu)$ and so the sum of cardinals each $<\mu$ indexed by ordinals $<\beta$ will be less than $\mu$.
@@
The specification(s) of a strong $\beta$-support exactly determine the orbit in the action of $\beta$-allowable permutations on supports to which it belongs: if two $\beta$-supports have the same specification, they are in the same orbit.
@@
It is straightforward to see that if $S$ is a $\beta$-support and if $\pi$ is a $\beta$-allowable permutation, and $S^{*}$ is the specification for $S$, that $S^{*}$ is also the specification for $\pi[S]$. The relationships between items in the support recorded in the specification are invariant under application of allowable permutations.
It remains to show that if $S$ and $T$ are supports, and $S^{*}=T^{*}$ is a specification for both, there is an allowable permutation $\pi$ such that $\pi[S]=T$.
We construct $\pi$ using the Freedom of Action Theorem (4.18). Define the $\beta$-approximation $\psi$ in the following way.
If $S_{\epsilon}=(M,A)$ for $M$ a near-litter, then $T_{\epsilon}=(N,A)$ for $N$ a near-litter, and we set $\psi_{A}(M^{\circ})=N^{\circ}$. Note that if $S_{\delta}=(M,A)$ for $\delta\neq\epsilon$, the fact that $S$ and $T$ have the same specification ensures that $T_{\delta}=(N,A)$.
If we have $S_{\epsilon}=(\{x\},A)$, we will have $T_{\epsilon}=(\{y\},A)$ for some $y$, and we will set $\psi_{A}(x)=y$. Again, if $S_{\delta}=(\{x\},A)$, then $T_{\delta}=(\{y\},A)$.
We then complete orbits of atoms, with the proviso that if $x\in M$ where $S_{\delta}=(M,A)$, then $\psi_{A}(x)\in N$ where $T_{\delta}=(N,A)$. During this process, we ensure that for every near-litter $M$ with $(M,A)\in{\tt rng}(S)$, we have $M\Delta M^{\circ}\subseteq{\tt dom}(\psi_{A})$, and similarly, whenever $(N,A)\in{\tt rng}(S)$, we have $N\Delta N^{\circ}\subseteq{\tt dom}(\psi_{A})^{-1}$. This process makes $\psi$ into an approximation, and ensures that $\psi_{A}^{*}(M)=N$.
We now need to check that $\psi$ is coherent. If $M$ is $A$-flexible and we assigned $\psi_{A}(M^{\circ})=N^{\circ}$, then $(M,A)=S_{\epsilon}$ and $(N,A)=T_{\epsilon}$ for some $\epsilon$. As $S$ and $T$ have the same specification, $M$ is $A$-flexible if and only if $N$ is $A$-flexible.
Now suppose that $M^{\circ}=f_{\delta,\gamma}(x,S^{\prime})$ where $\gamma={\tt min}(A_{1})$ and $\delta<{\tt min}(A_{2})$. Again, as $S$ and $T$ have the same specification, we have
  1. 1.
    $N^{\circ}=f_{\delta,\gamma}(y,T^{\prime})$;
  2. 2.
    there is a function $F$ such that $S_{F(\delta)}=({S^{\prime}}^{\uparrow A_{2}})_{\delta}$ and $T_{F(\delta)}=({T^{\prime}}^{\uparrow A_{2}})_{\delta}$; and
  3. 3.
    $\chi_{x,S^{\prime}}=\chi_{y,T^{\prime}}$.
As $\chi_{x,S^{\prime}}=\chi_{y,T^{\prime}}$, there is a $\delta$-allowable permutation $\pi$ such that $\pi[S^{\prime}]=T^{\prime}$ and $\pi(x)=y$. By the second property and the definition of $\psi$, we have $(\psi_{A_{2}})_{\delta}^{*}[S^{\prime}]=T^{\prime}=\pi[S^{\prime}]$. It remains to check that
$\psi_{A}^{+}``f_{\delta,\gamma}(x,S^{\prime})\sim f_{\delta,\gamma}(\pi(x),\pi^{\prime}[S])$
but this follows directly from the fact that $M^{\circ}=f_{\delta,\gamma}(x,S^{\prime})$ and $N^{\circ}=f_{\delta,\gamma}(\pi(x),\pi^{\prime}[S])$.
Finally, by the Freedom of Action Theorem (4.18), $\psi$ approximates some $\beta$-allowable permutation $\pi$, and in particular, $\pi$ satisfies $\pi[S]=T$, as required.
Proposition 4.30.
Under the inductive hypothesis that for each $\beta<\alpha$ we have $<\mu$ $\beta$-coding functions, there are less than $\mu$ orbits in supports under $\beta$-allowable permutations.
@@
Since the $\beta$-specifications [$\beta\leq\alpha$] precisely determine the orbits in strong supports under $\beta$-allowable permutations, and there are $<\mu$ $\beta$-specifications (on stated hypotheses) there are $<\mu$ such orbits.
Notice that we can give a kind of specification for any support $S$: give the specification for a strong support $T$ of which $S$ is an end extension and the index at which $S$ starts: this will determine the orbit in which $S$ lies in the allowable permutations. This establishes that the collection of orbits in the $\beta$-supports is no larger than the collection of orbits in the strong $\beta$-supports. These weak specifications are not unique. ∎
@@
There are less than $\mu$ coding functions for type 0 and for type $-1$
@@
We describe all coding functions for type 0. The orbit of a 0-support in the allowable permutations is determined by the positions in the support occupied by near-litters, and for each position in the support occupied by a singleton, the positions, if any, of the near-litters in the support which include it. There are no more than $2^{\kappa}$ ways to specify an orbit. Now for each such equivalence class, there is a natural partition of type $-1$ into near-litters, singletons, and a large complement set. The partition has $\nu<\kappa$ elements, and there will be $2^{\nu}\leq 2^{\kappa}$ coding functions for that orbit in the supports, determined by specifying for each compartment in the partition whether it is to be included or excluded from the set computed from a support in that orbit. So there are no more than $2^{\kappa}<\mu$ coding functions over type 0.
Lemma 4.32.
Under the inductive hypothesis that for each $\beta<\alpha$ we have $<\mu$ $\beta$-coding functions, there are less than $\mu$ coding functions for $\alpha$.
Proof.
By lemma 4.31 we may assume $\alpha$ is positive. Note that we already know that there are $<\mu$ $\alpha$-specifications, on the stated inductive hypothesis (proposition 4.28).
We specify an object $X\in\tau_{\alpha}$ and an $\alpha$-support $S$ for $X$, and develop a recipe for the coding function $\chi_{X,S}$ which can be used to see that there are $<\mu$ $\alpha$-coding functions (assuming of course that we know that things worked out correctly for $\beta<\alpha$).
$X=B_{\alpha}$, where $B$ is a subset of $\tau_{\beta}$. ($\beta<\alpha$ is chosen arbitrarily here, and could for example be chosen to be 0).
We define $S_{b}$ as the designated strong support for $b$, and $T_{b}$ as the canonical extension to a strong support of $S_{b}^{\uparrow\alpha}+S$.
For each $b\in B$ there is a support $S_{b}$ chosen as above, from which the support $T_{b}$ can be computed as described above. If $b^{\prime}\in B$ is in the range of the same coding function $\chi_{b,S_{b}}$ as $b$, $S_{b^{\prime}}$ is $\pi[S_{b}]$ for some $\beta$-allowable $\pi$ with $\pi(b)=b^{\prime}$. If we have the further condition that $T_{b}$ and $T_{b^{\prime}}$ have the same specification, it follows that there is a permutation $\pi_{2}$ such that $\pi_{2}[T_{b}]=T_{b^{\prime}}$. Note that $(\pi_{2})_{\beta}[S_{b}]=S_{b^{\prime}}$, from which it follows that $\pi^{-1}\circ(\pi_{2})_{\beta}$ fixes $b$, since it fixes all elements of $S_{b}$, so $b^{\prime}=\pi(b)=(\pi_{2})_{\beta}(b)$, from which it follows that $\pi_{2}(\{b\}_{\alpha})=\{b^{\prime}\}_{\alpha}$ so $\{b\}_{\alpha}$ and $\{b^{\prime}\}_{\alpha}$ are in the range of the same coding function $\chi_{\{b\}_{\alpha},T_{b}}$. Now there are $<\mu$ possible specifications of a coding function $\chi_{b,S_{b}}$ [we know that there are $<\mu$ $\beta$-coding functions] followed by a specification for $T_{b}$ [we know that there are $<\mu$ $\alpha$-specifications], so by this procedure we describe a family of $<\mu$ coding functions $\chi_{\{b\}_{\alpha},T_{b}}$ whose range covers all type $\alpha$ singletons of elements of $B$.
We claim that $\chi_{X,S}$ can be defined in terms of the orbit of $S$ in the allowable permutations and the set of coding functions $\chi_{\{b\}_{\alpha},T_{b}}$ for $b\in B$. There are $<\mu$ coding functions $\chi_{\{b\}_{\alpha},T_{b}}$ for $b\in\tau_{\beta}$, and so there are $<\mu$ sets of coding functions of this kind, because $\mu$ is strong limit, and we have shown above in proposition 4.30 that there are $<\mu$ orbits in the strong $\alpha$-supports under allowable permutations, so this will imply that there are $<\mu$ $\alpha$-coding functions, which will further imply that there are $\leq\mu$ elements of type $\alpha$ (it is obvious that there are $\geq\mu$ elements of each type).
The definition that we claim works is that $\chi_{X,S}(U)=B^{\prime}_{\alpha}$, where $B^{\prime}$ is the set of all $\bigcup(\chi_{\{b\}_{\alpha},T_{b}}(U^{\prime})\cap\pi_{\beta}$) for $b\in B$ and $U$ a terminal segment of $U^{\prime}$. Clearly this definition depends only on the orbit of $S$ and the set of coding functions $\chi_{\{b\}_{\alpha},T_{b}}$ derived from $B$ as described above. Before we know that this is actually the coding function desired, we will write it as $\chi_{X,S}^{*}$.
The function we have defined is certainly a coding function, in the sense that $\chi_{X,S}^{*}(\pi[U])=\pi(\chi_{X,S}^{*}(U))$. What requires work is to show that $\chi_{X,S}^{*}(S)=X$, from which it follows that it is in fact the intended function.
Clearly each $b\in B$ belongs to $\chi^{*}_{X,S}(S)$ as defined, because
$b=\bigcup(\chi_{\{b\}_{\alpha},T_{b}}(T_{b})\cap\tau_{\beta})$, and $T_{b}$ has $S$ as a terminal segment.
An arbitrary $c\in\chi_{X,S}^{*}(S)$ is of the form $\bigcup(\chi_{\{b\}_{\alpha},T_{b}}(U)\cap\tau_{\beta})$, where $U$ has $S$ as a terminal segment and of course must be in the orbit of $T_{b}$ under allowable permutations, so some $\pi_{0}[T_{b}]=U$. Now observe that $\pi_{0}[S]=S$, so $\pi_{0}(X)=X$, so $(\pi_{0})_{\beta}``B=B$. Further $(\pi_{0})_{\beta}(b)=c$, so in fact $c\in B$ which completes the argument. The assertion $(\pi_{0})_{\beta}(b)=c$ might be thought to require verification: the thing to observe is that
$c=\bigcup(\chi_{\{b\}_{\alpha},T_{b}}(U)\cap\tau_{\beta})=\bigcup(\pi_{0}(\chi_{\{b\}_{\alpha},T_{b}}(T_{b}))\cap\tau_{\beta})=\bigcup(\pi_{0}(\{b\}_{\alpha})\cap\tau_{\beta})$
=($=\bigcup(\{(\pi_{0})_{\beta}(b)\}_{\alpha}\cap\tau_{\beta})=(\pi_{0})_{\beta}(b).$
@@ 4.3 The structure is a model of predicative TTT
Suppose that we are considering the existence of $\{x:\phi^{s}\}$, where $\phi$ is a formula of the language of TST with $\in$ translated as $E$, and $s$ is a strictly increasing sequence of types. The truth value of each subformula of $\phi$ will be preserved if we replace each $u$ of type $s(i)$ with $\pi_{A_{s,i}}(u)$, where $A_{s,i}$ is the set of all $s_{k}$ for $i\leq k\leq j+1$ [$x$ being of type $s(j)$, and there being no variables of type higher than $s(j+1)$]: $\pi_{A_{s,i}}(x)E\pi_{A_{s,i+1}}(y)$ is equivalent to $(\pi_{A_{s,i+1}})_{s(i)}(x)E\pi_{A_{s,i+1}}(y)$, which is equivalent to $xEy$ by the observation above. The formula $\phi$ will contain various parameters $a_{i}$ of types $s(n_{i})$ and it is then evident that the set $\{x:\phi^{s}\}$ will be fixed by any $s(j+1)$-allowable permutation $\pi$ such that $\pi_{A_{s,n_{i}}}$ fixes $a_{i}$ for each $i$. But this means that $\{x:\phi^{s}\}_{s(j+1)}$ is symmetric and belongs to type $s(j+1)$: we can merge the supports of the $a_{i}$’s (with suitable raising of indices) into a single $s(j+1)$-support. Notice that we assumed the predicativity condition that no variable more than one type higher than $x$ appears (in the sense of TST).
@@
It should be noted that the construction given here is in a sense a Frankel-Mostowski construction, though we have no real need to reference the usual FM constructions in ZFA here. Constructions analogous to Frankel-Mostowski constructions can be carried out in TST using permutations of type 0; here we are doing something much more complicated involving many permutations of type $-1$ which intermesh in precisely the right way. Our explanation of our technique is self-contained, but we do acknowledge this intellectual debt.

### v23

@@ A $\beta$-support $S$ is called strong if it satisfies the additional properties that
The elements of the range of a $\beta$-specification are built from finite sets of type indices below $\beta$, sets of (pairs of) elements of $\kappa$ (support domain elements), and $\gamma$-coding functions for $\gamma<\beta$. As $\beta<\lambda\leq\mu$, there are less than $\mu$ possible finite sets of type indices below $\beta$. Since $\beta<\lambda\leq{\tt cf}(\mu)$, the sum of cardinals each $<\mu$ indexed by ordinals $<\beta$ will be less than $\mu$, so there are less than $\mu$ possible coding functions that can be used here. Hence the elements of the range of such a specification are taken from a set of size less than $\mu$, and so the cardinality of the set of specifications is bounded by $\nu^{<\kappa}$ where $\nu<\mu$. Now, $\nu^{<\kappa}$ is the cardinality of the set of functions from some small ordinal to $\nu$, which is less than or equal to the cardinality of the power set of $\nu\times\kappa$, which in turn is less than $\mu$ because $\mu$ is strong limit. ∎
Proposition 4.32.
Suppose that we already know that there are $<\mu$ $\gamma$-coding functions for each $\gamma<\beta$. Then there are less than $\mu$ orbits in supports under $\beta$-allowable permutations.
Proof.
For any support $S$ that is not necessarily strong, there is a suitable weaker notion of specification: give the specification for a strong support $T$ of which $S$ is an initial segment, together with the domain of $S$. These weak specifications are not unique, but they do determine the orbit in which $S$ lies in the allowable permutations. Since there are less than $\mu$ specifications by proposition 4.31, and $\kappa<\mu$, we conclude that there are less than $\mu$ orbits in the $\beta$-supports under the allowable permutations. ∎
4.4 Types are of size $\mu$ (because there are $<\mu$ coding functions per type)
We will prove this cardinality bound on coding functions by induction on $\beta\leq\alpha$. It is important to note that this inductive proof occurs internally to a single instance of the main inductive step; that is, $\alpha$ remains fixed, and we do not need to record the fact that the bound on coding functions holds in our main list of inductive hypotheses (I1)–(I11).
This argument will proceed by showing that all coding functions can be written in a particular form, in terms of coding functions at some lower level $-1\neq\gamma<\beta$. In particular, this result will depend on the existence of some ordinal $\gamma\in\beta$, so we prove the result for types 0 and $-1$ explicitly here first.
@@ 4.4 Types are of size $\mu$ (because there are $<\mu$ coding functions per type)
There are less than $\mu$ coding functions for type 0 and for type $-1$.
@@
We describe all coding functions for type 0. The orbit of a strong 0-support in the allowable permutations is determined by the positions in the support occupied by near-litters, and for each position in the support occupied by a singleton, the positions, if any, of the near-litters in the support which include it, and for each position in the support occupied by a near-litter, the positions, if any, occupied by near-litters which are near it (this is exactly the information contained in the 0-specification of the support). There are no more than $2^{\kappa}$ ways to specify an orbit. Now for each such support, there is a natural partition of type $-1$ into near-litters, singletons, and a large complement set. The partition has $\nu<\kappa$ elements, and there are $2^{\nu}\leq 2^{\kappa}$ objects of type 0 with this support, each determined by specifying for each compartment in the partition whether it is to be included or excluded from the set. So there are $2^{\nu}\leq 2^{\kappa}$ coding functions for that orbit in the supports, and so for any support (that need not be strong), there are at most $2^{\kappa}$ coding functions for its orbit, by first extending it to a strong support. So there are no more than $2^{\kappa}<\mu$ coding functions over type 0.
With those two cases handled, we return to focus on the case where there is some ordinal $\gamma\in\beta$.
Definition 4.34 (combination).
Let $\gamma\in\beta$ be ordinals at most $\alpha$. Let $s$ be a set of $\beta$-coding functions. We define the combination of $s$ along $\gamma$ to be the function $C_{s,\gamma}$ from $\beta$-supports to $\tau_{\beta}$ given by
$C_{s,\gamma}(S)=\left(\bigcup_{\chi\in s}\bigcup_{T\supseteq S}\chi(T)\cap\tau_{\gamma}\right)_{\beta}.$
That is, for a support $S$, the $\gamma$-extension of the model element $C_{s,\gamma}(S)$ is precisely the union of the $\gamma$-extensions of model elements obtained by applying coding functions from $s$ to supports extending $S$.
That $C_{s,\gamma}(S)$ is always a model element requires proof: by symmetric comprehension (proposition 3.29), it suffices to show that when a $\beta$-allowable permutation $\pi$ fixes $S$, $\pi_{\gamma}$ fixes the set
$\bigcup_{\chi\in s}\bigcup_{T\supseteq S}\chi(T)\cap\tau_{\gamma},$
but this follows from the fact that for any $\beta$-allowable $\pi$,
$\pi_{\gamma}``\left(\bigcup_{\chi\in s}\bigcup_{T\supseteq S}\chi(T)\cap\tau_{\gamma}\right)$ $=\bigcup_{\chi\in s}\bigcup_{T\supseteq S}\pi_{\gamma}``(\chi(T)\cap\tau_{\gamma})$
$=\bigcup_{\chi\in s}\bigcup_{T\supseteq S}\pi(\chi(T))\cap\tau_{\gamma}$
$=\bigcup_{\chi\in s}\bigcup_{T\supseteq S}\chi(\pi[T])\cap\tau_{\gamma}$
$=\bigcup_{\chi\in s}\bigcup_{T\supseteq\pi[S]}\chi(T)\cap\tau_{\gamma}.$
The same calculation shows that for any orbit of $\beta$-supports $o$ under the action of allowable permutations, the restriction $C_{s,\gamma}\restriction o$ is a coding function.
Our claim is that every $\beta$-coding function can be written in this form, where the elements of the set $s$ are particular coding functions of singletons of objects of type $\gamma$. We will make the reference of “particular” here more precise.
Definition 4.35 (raised singleton).
Let $\gamma\in\beta$. A $\beta$-coding function $\chi$ is called a raised $\gamma$-singleton if there is a $\beta$-support $S$ and $(x,U)\in\tau_{\gamma}^{+}$ such that $\chi=\chi_{\{x\}_{\beta},S+U^{\uparrow\beta}}$.
Lemma 4.36.
Let $\gamma\in\beta$ be ordinals at most $\alpha$, and let $\chi$ be a $\beta$-coding function. Then there is a set $s$ of raised $\gamma$-singletons such that $\chi=C_{s,\gamma}\restriction{\tt dom}(\chi)$.
@@
Let $\chi=\chi_{x,S}$. The set we claim works is
$s=\{\chi_{\{y\}_{\beta},S+U^{\uparrow\beta}}:y\in x\cap\tau_{\gamma}\wedge U\text{ supports }y\}.$
Since $C_{s,\gamma}\restriction{\tt dom}(\chi)$ is a coding function, it suffices to check that $C_{s,\gamma}(S)=x$, which we show by $\gamma$-extensionality.
If $y\in C_{s,\gamma}(S)\cap\tau_{\gamma}$, then there is some $z\in x\cap\tau_{\gamma}$, some support $U$ of $z$, and $T\supseteq S$ such that
$y\in\chi_{\{z\}_{\beta},S+U^{\uparrow\beta}}(T).$
Hence there is a $\beta$-allowable $\pi$ such that $\pi[S+U^{\uparrow\beta}]=T$; in particular, $\pi[S]=S$ as $T$ is an extension of $S$, so $\pi(x)=x$. Therefore,
$\pi_{\gamma}^{-1}(y)\in\chi_{\{z\}_{\beta},S+U^{\uparrow\beta}}(\pi^{-1}[T])=\chi_{\{z\}_{\beta},S+U^{\uparrow\beta}}(S+U^{\uparrow\beta})=\{z\}_{\beta},$
giving $\pi_{\gamma}^{-1}(y)\in x$ and so $y\in x$ as required.
Conversely, if $y\in x\cap\tau_{\gamma}$, then for any support $U$ of $y$, we immediately obtain $y\in C_{s,\gamma}(S)$ from the fact that $\chi_{\{y\}_{\beta},S+U^{\uparrow\beta}}\in s$. ∎
@@
There are less than $\mu$ coding functions for every type index $\beta\leq\alpha$.
@@
We may assume by induction that there are less than $\mu$ coding functions of every type $\gamma<\beta$. Since the cases for $\beta=0,-1$ were proven in lemma 4.33, we may additionally assume that there exists some ordinal $\gamma\in\beta$. By lemma 4.36, every $\beta$-coding function can be written in the form $C_{s,\gamma}\restriction d$, where $s$ is a set of raised $\gamma$-singletons and $d$ is an orbit of $\beta$-supports under the action of allowable permutations. By proposition 4.32, there are less than $\mu$ such orbits, so it suffices to show there are less than $\mu$ possible sets of raised $\gamma$-singletons; indeed, as $\mu$ is a strong limit, it further suffices to simply show that there are less than $\mu$ raised $\gamma$-singletons.
We show that any raised singleton $\chi_{\{x\}_{\beta},S+U^{\uparrow\beta}}$ is determined by ${\tt dom}(S)$, the orbit of $S+U^{\uparrow\beta}$ under the allowable permutations, and the $\gamma$-coding function $\chi_{x,U}$. Indeed, suppose that $S$ and $T$ have the same domain, $S+U^{\uparrow\beta}$ and $T+V^{\uparrow\beta}$ share the same orbit under the $\beta$-allowable permutations, and the coding functions $\chi_{x,U}$ and $\chi_{y,V}$ coincide. Then there is a $\beta$-allowable $\pi$ such that $\pi[S]=T$ and $\pi_{\gamma}[U]=V$, and additionally there is a $\gamma$-allowable $\pi^{\prime}$ such that $\pi^{\prime}[U]=V$ and $\pi^{\prime}(x)=y$. Therefore, we additionally have that $\pi_{\gamma}(x)=y$ as $U$ supports $x$ and both $\pi_{\gamma}$ and $\pi^{\prime}$ have the same action on $U$. Hence, the two raised singletons must coincide.
It now simply remains to note that there are less than $\mu$ such orbits of $\beta$-supports under the allowable permutations by proposition 4.32, and that we assume by inductive hypothesis that there are less than $\mu$ coding functions of type $\gamma$, in order to conclude the desired result. ∎
@@ 4.5 The structure is a model of predicative TTT
Suppose that we are considering the existence of $\{x:\phi^{s}\}$, where $\phi$ is a formula of the language of TST with $\in$ translated as $E$, and $s$ is a strictly increasing sequence of types [^12] We use notation $s(i)$ instead of $s_{i}$ to avoid undue nesting of subscripts, in what follows.. Note that the TTT type of $x$ is some $s(j)$ and the TTT type of $\{x:\phi^{s}\}$ is $s(j+1)$. For any $s(j+1)$-allowable permutation $\pi$, the truth value of each subformula of $\phi$ will be preserved if we replace each $u$ of type $s(i)$ with $\pi_{A_{s,i}}(u)$, where $A_{s,i}$ is the set of all $s(k)$ for $i\leq k\leq j+1$ [there being no variables of type higher than $s(j+1)$]: $\pi_{A_{s,i}}(x)E\pi_{A_{s,i+1}}(y)$ is equivalent to $(\pi_{A_{s,i+1}})_{s(i)}(x)E\pi_{A_{s,i+1}}(y)$, which is equivalent to $xEy$ by the observation above. The formula $\phi$ will contain various parameters $a_{i}$ of types $s(n_{i})$ and it is then evident that the set $\{x:\phi^{s}\}$ will be fixed by any $s(j+1)$-allowable permutation $\pi$ such that $\pi_{A_{s,n_{i}}}$ fixes $a_{i}$ for each $i$. But this means that $\{x:\phi^{s}\}_{s(j+1)}$ is symmetric and belongs to type $s(j+1)$: we can merge the supports of the $a_{i}$’s (with suitable raising of indices) into a single $s(j+1)$-support. Notice that we assumed the predicativity condition that no variable more than one type higher than $x$ appears (in the sense of TST).
@@
It should be noted that the construction given here is in a sense a Fraenkel–Mostowski construction, though we have no real need to reference the usual FM constructions in ZFA here. Constructions analogous to Fraenkel–Mostowski constructions can be carried out in TST using permutations of type 0; here we are doing something much more complicated involving many permutations of type $-1$ which intermesh in precisely the right way. Our explanation of our technique is self-contained, but we do acknowledge this intellectual debt.